OpenStudy (anonymous):

The volume of a rectangular slab of concrete needs to exceed 72 cu ft. If the lenght is 12ft and the width is 8 ft how thick does the concrete slab need to be

OpenStudy (campbell_st):

\[V = l \times w \times h\] in this question you have \[72 = 12 \times 8 \times h\] \[h = \frac{72}{96}\] I'll let you evaluate th height... it will be a decimal... and the meaurement is feet